University of Wisconsin-Madison

South Pole Weekly Report: Week ending January 9

Population

South Pole Station Population is currently at 223, a decrease of 16 from the previous week. Three IceCube people are stuck in McMurdo for the weekend, so the IceCube population is unchanged from last week, 38 against a plan of 41. One of our winter-overs left for a recreation break in New Zealand, and the other is scheduled for a break later this month.

The B2 Science area has taken on a new feeling as staff focus on commissioning, data acquisition integration, and other science support activities.

The on-ice population will see a dramatic change in the coming week as the drillers depart. Thirty departures are scheduled and five arrivals, including those stuck in McMurdo.

Drilling and logistics

It has been a busy week at the drill camp. The drilling and deployment towers, main heating plant, Rodriguez well System, pre-heater system, main hose reel, and the optical module lab have all been placed in storage for the winter.

Remaining machinery are being readied for storage. Complete inventory of drill, plumbing and electrical parts began on Friday and is scheduled to be done next week. The seasonal equipment site has dwindled to a few buildings.

We shipped approximately 3,500 pounds of do not freeze equipment last week. IceCube Lab and Enhanced Hot Water Drill components were included. We also coordinated with McMurdo logistics staff to ship and store seven spools of drill hose.

With drill teams packing up equipment and buildings, only a few buildings in the seasonal equipment site remain in use.
With drill teams packing up equipment and buildings, only a few buildings in the seasonal equipment site remain in use.

Engineering Support and Commissioning

During commissioning of string 14 it was discovered that several Digital Optical Modules were unconnected. A small nick in the surface and down-hole cables in the IceTop pit was discovered, and we are working on repairing them before engineering support personnel leave the pole on January 12. If a permanent repair cannot be accomplished in this time frame, a plywood vault will be constructed around the damaged area, allowing easier access for repair next season.

Final preparations for the cables into the IceCube Lab were completed. A special thanks to those that worked on this complicated cable operation.

In-ice commissioning activities began on Friday. A couple of strings have issues, including cable damage. We are investigating the issues and working on integrating the Digital Optical Modules into the detector array.

IceTop

The IceTop headquarters, also known as the "Purple Palace," was emptied and disassembled. Remaining items were sent for scrap for prepared for return to the University of Delaware or the UW-Madison Physical Sciences Laboratory.

All IceTop tanks are continuing to freeze and working as planned. Data transfer is being monitored since delays have been noticed.